# Tony Lindeberg

### v3

@@ 3.2 Time-causal bandpass wavelet representations
where the difference-of-the-time-causal-limit kernel $\Delta\Psi$ is given by
$\Delta\Psi(t;\;\tau_{k},\tau_{k-1},c)=\Psi(t;\tau_{k})-\Psi(t;\tau_{k-1},c)$ (99)
for
@@
Specifically, if such a representation is computed between a finest minimum temporal scale level $\tau_{1}$, for the specific choice of defining $L(t;\tau_{0},c)=f(t)$ in the boundary case, and a coarsest maximum temporal scale level $\tau_{K}$, then reconstruction of the original signal or the representation at any other temporal scale level is trivial from such a representation, by merely adding the corresponding bandpass representations over the temporal scale channels for all the temporal scale levels:
@@ 3.3 Bandpass representations based on the second-order derivatives of the Gaussian kernel
Thus, we have that, up to a scalar normalization factor and a normalization with respect to the scale parameter, the Ricker wavelet, also known as the Mexican hat wavelet, and which approximately encodes differences in a Gaussian-based-scale-space representation between temporal scales.
@@
Specifically, regarding the second-order derivative of the non-causal Gaussian kernel, it also follows from the diffusion equation interpretation of the non-causal 1-D scale-space representation (104), that provided that second-order Gaussian-derivative-based responses are computed over a continuum of scale levels over some range $t_{1},t_{2}]$, then the signal at the finer scale level $t_{2}$ from the signal at the coarser scale level $t_{2}$ in combination with the second-order derivative responses over this scale range according to
@@
This expression can be seen as a non-causal analogue of the corresponding reconstruction (101) from a time-causal bandpass representation, again demonstrating a close structural relationship to regard the temporal derivatives of the time-causal limit kernels as time-causal analogues of the derivatives of the regular Gaussian kernel.
@@
Regarding the difference-of-the-time-causal-limit kernel (99) used in the temporal bandpass wavelets representation in Section 3.2, it follows from the differential equation interpretation (30) in terms of first-order integrators coupled in cascade that this difference kernel does instead approximate the first-order temporal derivative of the time-causal limit kernel according to
$\Delta\Psi(t;\;\tau_{k},\tau_{k-1},c)\approx-\mu_{k}\,\Psi_{t}(t;\;\tau_{k},c),$ (112)
and also with the notable difference that the difference between the temporal scale levels $\tau_{k}$ and $\tau_{k-1}$ has to be macroscopic of the form
@@ 4.1 Discretization of the first-order integrators
where we here assume that the discrete signals in Equation (115) are sampled with a grid spacing $\Delta t$ relative to the continuous signals in Equation (114). Then, it can be shown that (Lindeberg 2025 Appendix B.3) the discrete recurrence relation (115) constitutes a true discretization of the continuous differential equation (114) provided that
@@ 4.2 Classification of discrete temporal smoothing kernels
This result also builds upon a classification of discrete variation-diminishing convolution transformations by Schoenberg (1948), which states that a discrete kernel guarantees non-creation of new zero-crossings in the convolved signal if and only if its generating function is of the form
@@ 5 Time-causal wavelet representations for discrete signals
for some suitably selected value of the scale normalization parameter $\gamma>0$, we thereby obtain truncated numerical approximations of the discrete analogue $L_{\zeta^{n}}\colon{\mathbb{Z}}\times{\mathbb{Z}}\rightarrow{\mathbb{R}}$ of the temporal derivative of order $n$ of the time-causal scale-space representation of the signal $f$ according to
@@ 6 Characterizations of scaling properties in discretizations of the continuous theory
The theory for time-causal and time-recursive wavelets in Section 2 is developed for fully continuous signals. In this section, we will characterize how well the continuous scaling properties carry over to a discrete implementation, as based on the methodology proposed in Sections 4 and 5, and depending on the relationship between the temporal scale levels in relation to the sampling distance in the signal.
@@ 6.2 Scale selection properties for local blob-like and edge-like structures in the signal
Such points computed in a temporal scale-space representation are referred to as scale-space extrema, and constitute a generalization of a corresponding methodology developed for a spatial scale-space representation (Lindeberg 1998). Specifically, based on the theory described in Section 2.7.2, in case the time-causal limit kernel would be replaced by a non-causal Gaussian kernel, this methodology would in the ideal continuous case determine the inherent temporal duration of the corresponding idealized temporal blob and edge models according to (57) and (62).
Figure 6 shows graphs of the dependency of the resulting scale estimates for a range of reference scale levels $\tau_{\text{ref}}$, when detecting scale-space extrema in this way for the reference blob and edge signals $f_{\text{blob}}(t;\;\tau_{\text{ref}})$ and $f_{\text{edge}}(t;\;\tau_{\text{ref}}$ according to (135) and (136) when using time-causal wavelet representations for a fixed reference scale $\tau_{0}=1$ in (12) using the distribution parameter set to either $c=2$ or $c=\sqrt{2}$. Given the discrete nature of both the temporal scale levels and the temporal sampling, we used local parabolic interpolation around each extremum over temporal scales to increase the accuracy in the resulting estimates of the temporal scales $\hat{\tau}$.
@@ 6.3 Temporal delays of the discrete approximations of the time-causal limit kernel
Figure 7 shows the result of computing these characterizations of the temporal delay for the distribution parameter $c\in\{\sqrt{2},2\}$, and with the estimate of $t_{\text{max}}$ interpolated to subquantization resolution using parabolic interpolation. For comparison, the graphs also shows the dependencies on the temporal scale levels for the continuous temporal delay estimates according to Equations (28) and (29). As can be seen from the graphs, the measures of the temporal delays are significantly shorter for the distribution parameter $c=2$ compared to the distribution parameter $c=\sqrt{2}$. We can also note that the measures of the temporal delays are shorter for the discrete approximations of the time-causal limit kernel compared to the continuous time-causal limit kernel.
@@
Figure 10: The result of computing discrete approximations of convolutions with the temporal derivatives of the time-causal limit kernel $L_{\zeta^{n}}$ for orders $n=1$ and $n=2$ from an auditory signal with a percussion sound from a cymbal, here over the scale range $\sigma=\sqrt{\tau}\in[1/8,1024]$ and using the scale normalization power $\gamma=1$ corresponding to $L_{1}$-normalization over scales of the underlying temporal derivatives of the time-causal limit kernel. In the bottom row, we have also shown the result of computing the quasi quadrature measure ${\cal Q}L=\sqrt{L_{\zeta}^{2}+C\,L_{\zeta\zeta}^{2}}$ for the relative weighting parameter $C=1/\sqrt{2}$. (Horizontal axes: time $\in[0,150000]$ relative to a sampling frequency of 22.05 kHz. Vertical axes in the time-causal wavelet representations: Effective scale = $\log_{2}\sqrt{\tau}$.) (In the time-causal wavelet representations, red denotes positive values and blue denotes negative values.) (The values on the vertical axis of the signal have been shifted to make the DC level approximately zero.)
@@ 7 Experiments
Let us also illustrate the effects of applying the time-causal frequency analysis to two realistic signals: Figure 9 shows such results for an electrocardiography signal from the PhysioNet service (http://www.physionet.org) from the MIT-BIH Arrhythmia database (Plawiak 2017). As can be seen from the results, there are strong values in the wavelet transforms near the transients in the signal. At coarser temporal scales, the temporal delays become longer. The temporal delays are, however, signficantly shorter when the distribution parameter is $c=2$ compared to using $c=\sqrt{2}$.
Figure 9 shows such results for an auditory percussion sound of a cymbal. As we can see, the time-causal wavelet transform largely responds over three ranges of temporal scales, with different durations of the responses over these ranges of temporal scales. In this way, the proposed time-causal wavelet transform leads to an intuitive separation of the temporal variations in the signal over different temporal scales.
@@ 8 Summary and discussion
Then, we have in Section 4 turned to the problem of numerically approximating the time-causal limit kernel and its temporal derivatives for discrete signals. In this regard, first-order recursive filters constitute the natural analogue to the first-order integrators that equivalently describe the temporal smoothing with truncated exponential kernels, and which constitute the main computational primitives in the continuous theory. Thereby, we have describe a canonical way to compute discrete time-causal wavelet representations in Section 5, and also shown how this representation relates to bandpass representations, as can also be defined from continuous wavelets. Notably, the proposed implementation method is fully time-recursive, and does therefore lend itself directly for real-time processing on regular signal processing architectures.
@@
Based on these results, and the abovementioned principled theoretical foundations, we propose that this notion of time-causal wavelet analysis could be a valuable tool for signal processing tasks, where streams of signals are to be processed in real time, specifically for signals that may contain local variations over a rich span of temporal scales, or more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@
Concerning further applications of the theory, we have intentionally in this presentation focused solely on theoretical and generic properties of the proposed families of time-causal wavelet representations, and not here aimed at developing more application-oriented experiments beyond the proof-of-concept studies in Sections 7 and 6.2. The motivation for this is that we want the results to be as timeless as possible, and not depending on the degree of match between the assumptions underlying the theory and properties of the experimental data, which may first of all be different for different application domains, and may furthermore warrant the inclusion of additional more specialized computational mechanisms as well as integration with other signal processing approaches to build a full-fledge application.
@@
For the closely related notion of a time-causal and time-recursive analogue of Gabor filtering, corresponding to a time-causal analogue of the Morlet wavelet (Morlet et al. 1982a; 1982b), see Lindeberg (2025). For more general notions of a non-causal wavelets based on Hermite transforms, and which constitute extensions of the wavelets based on first- and second-order derivatives of the Gaussian kernel to higher orders of differentiation, see Markett (1993), Glaeske (2000) and Pandey and Phukan (2020).

### v6

@@ 3.2 Time-causal bandpass wavelet representations
where the difference-of-the-time-causal-limit-kernels (DoT) kernel $\Delta\Psi$ is given by
$\Delta\Psi(t;\;\tau_{k},\tau_{k-1},c)=\Psi(t;\tau_{k},c)-\Psi(t;\tau_{k-1},c)$ (99)
for (Lindeberg 2016 Equation (18))
@@
Specifically, if such a representation is computed between a finest minimum temporal scale level $\tau_{1}$, for the specific choice of defining $L(t;\tau_{0},c)=f(t)$ in the boundary case, and a coarsest maximum temporal scale level $\tau_{K}$, then reconstruction of the original signal or the representation at any other temporal scale level is straightforward from such a representation, by merely adding the corresponding bandpass representations over the temporal scale channels for all the temporal scale levels:
@@ 3.3 Bandpass representations based on the second-order derivatives of the Gaussian kernel
Thus, we have that, up to a scalar normalization factor and a normalization with respect to the scale parameter, the Ricker wavelet, also known as the Mexican hat wavelet, approximately encodes differences in a Gaussian-based scale-space representation between temporal scales.
@@
Specifically, regarding the second-order derivative of the non-causal Gaussian kernel, it also follows from the diffusion equation interpretation of the non-causal 1-D scale-space representation (104), that provided that second-order Gaussian-derivative-based responses are computed over a continuum of scale levels over some range $t_{1},t_{2}]$, then the signal at the finer scale level $t_{2}$ can be reconstructed from the signal at the coarser scale level $t_{2}$ in combination with the second-order derivative responses over this scale range according to
@@
This expression can be seen as a non-causal analogue of the corresponding reconstruction (101) from a time-causal bandpass representation, again demonstrating a close structural relationship to regard the temporal derivatives of the time-causal limit kernel as time-causal analogues of the derivatives of the regular Gaussian kernel.
@@
Regarding the difference-of-the-time-causal-limit kernel (99) used in the temporal bandpass wavelets representation in Section 3.2, it follows from the differential equation interpretation (30) in terms of first-order integrators coupled in cascade, that this difference kernel does instead represent the first-order temporal derivative of the time-causal limit kernel according to
$\Delta\Psi(t;\;\tau_{k},\tau_{k-1},c)=-\mu_{k}\,\Psi_{t}(t;\;\tau_{k},c),$ (112)
and also with the notable difference that the difference between the temporal scale levels $\tau_{k}$ and $\tau_{k-1}$ now has to be macroscopic of the form
@@ 4.1 Discretization of the first-order integrators
where we here assume that the discrete signals in Equation (115) are sampled with a grid spacing $\Delta t$ relative to the continuous signals in Equation (114). Then, it can be shown that (Lindeberg 2025b Appendix B.3) the discrete recurrence relation (115) constitutes a true discretization of the continuous differential equation (114) provided that
@@ 4.2 Classification of discrete temporal smoothing kernels
The result that the time-causal smoothing operation should be discretized on the form (115) also builds upon a classification of discrete variation-diminishing convolution transformations by Schoenberg (1948), which states that a discrete kernel guarantees non-creation of new zero-crossings in the convolved signal if and only if its generating function is of the form
@@ 5 Time-causal wavelet representations for discrete signals
for some suitably selected value of the scale normalization parameter $\gamma>0$, we thereby obtain numerical approximations of the discrete analogue $L_{\zeta^{n}}\colon{\mathbb{Z}}\times{\mathbb{Z}}\rightarrow{\mathbb{R}}$ of the temporal derivative of order $n$ of the time-causal scale-space representation of the signal $f$ according to
@@ 6 Characterizations of scaling properties in discretizations of the continuous theory
The theory for scale-covariant time-causal and time-recursive wavelets in Section 2 is developed for fully continuous signals. In this section, we will characterize how well the continuous scaling properties carry over to a discrete implementation, as based on the methodology proposed in Sections 4 and 5, and depending on the relationship between the temporal scale levels in relation to the sampling distance in the signal.
@@ 6.2 Scale selection properties for local blob-like and edge-like structures in the signal
Such points computed in a temporal scale-space representation are referred to as scale-space extrema, and constitute a generalization of a corresponding methodology developed for a spatial scale-space representation (Lindeberg 1998a; 1998b). Specifically, based on the theory described in Section 2.7.2, in case the time-causal limit kernel would be replaced by a non-causal Gaussian kernel, this methodology would in the ideal continuous case determine the inherent temporal duration of the corresponding idealized temporal blob and edge models according to (57) and (62).
Figure 6 shows graphs of the dependency of the resulting scale estimates for a range of reference scale levels $\tau_{\text{ref}}$, when detecting scale-space extrema in this way for the reference blob and edge signals $f_{\text{blob}}(t;\;\tau_{\text{ref}})$ and $f_{\text{edge}}(t;\;\tau_{\text{ref}}$ according to (135) and (136), when using time-causal wavelet representations for a fixed reference scale $\tau_{0}=1$ in (12) using the distribution parameter set to either $c=2$ or $c=\sqrt{2}$. Given the discrete nature of both the temporal scale levels and the temporal sampling, we used local parabolic interpolation around each extremum over temporal scales to increase the accuracy in the resulting estimates of the temporal scales $\hat{\tau}$.
@@ 6.3 Temporal delays of the discrete approximations of the time-causal limit kernel
Figure 7 shows the result of computing these characterizations of the temporal delay for the distribution parameter $c\in\{\sqrt{2},2\}$, and with the estimate of $t_{\text{max}}$ interpolated to subquantization resolution using parabolic interpolation. For comparison, the graphs also show the dependencies on the temporal scale levels for the continuous temporal delay estimates according to Equations (28) and (29). As can be seen from the graphs, the measures of the temporal delays are significantly shorter for the distribution parameter $c=2$ compared to the distribution parameter $c=\sqrt{2}$. We can also note that the measures of the temporal delays are shorter for the discrete approximations of the time-causal limit kernel compared to the continuous time-causal limit kernel.
@@
Figure 10: The result of computing discrete approximations of convolutions with the temporal derivatives of the time-causal limit kernel $L_{\zeta^{n}}$ for orders $n=1$ and $n=2$ from a vibration signal, here over the scale range $\sigma=\sqrt{\tau}\in[1/8,1024]$ and using the scale normalization power $\gamma=1$ corresponding to $L_{1}$-normalization over scales of the underlying temporal derivatives of the time-causal limit kernel. In the bottom row, we have also shown the result of computing the quasi quadrature measure ${\cal Q}L=\sqrt{L_{\zeta}^{2}+C\,L_{\zeta\zeta}^{2}}$ for the relative weighting parameter $C=1/\sqrt{2}$. (Horizontal axes: time $\in[0,2000]$. Vertical axes in the time-causal wavelet representations: Effective scale = $\log_{2}\sqrt{\tau}$.) (In the time-causal wavelet representations, red denotes positive values and blue denotes negative values.) (The values on the vertical axis of the signal have been shifted to make the DC level approximately zero.)
Input signal
first-order derivative for $c=\sqrt{2}$ first-order derivative for $c=2$
Refer to caption Refer to caption
Refer to caption Refer to caption
quasi quadrature for $c=\sqrt{2}$ quasi quadrature for $c=2$
Refer to caption Refer to caption
Figure 11: The result of computing discrete approximations of convolutions with the temporal derivatives of the time-causal limit kernel $L_{\zeta^{n}}$ for orders $n=1$ and $n=2$ from an auditory signal with a percussion sound from a cymbal, here over the scale range $\sigma=\sqrt{\tau}\in[1/8,16384]$ and using the scale normalization power $\gamma=1$ corresponding to $L_{1}$-normalization over scales of the underlying temporal derivatives of the time-causal limit kernel. In the bottom row, we have also shown the result of computing the quasi quadrature measure ${\cal Q}L=\sqrt{L_{\zeta}^{2}+C\,L_{\zeta\zeta}^{2}}$ for the relative weighting parameter $C=1/\sqrt{2}$. (Horizontal axes: time $\in[0,150000]$ relative to a sampling frequency of 44.1 kHz. Vertical axes in the time-causal wavelet representations: Effective scale = $\log_{2}\sqrt{\tau}$.) (In the time-causal wavelet representations, red denotes positive values and blue denotes negative values.) (The values on the vertical axis of the signal have been shifted to make the DC level approximately zero.)
@@ 7 Experiments
Let us also illustrate the effects of applying the time-causal frequency analysis to three real-world signals: Figure 9 shows such results for an electrocardiography signal from the PhysioNet service (http://www.physionet.org) from the MIT-BIH Arrhythmia database, item “NSR/100m (0).mat” from (Plawiak 2017). As can be seen from the results, there are strong values in the wavelet transforms near the transients in the signal. At coarser temporal scales, the temporal delays become longer. The temporal delays are, however, significantly shorter when the distribution parameter is $c=2$ compared to using $c=\sqrt{2}$.
Figure 10 shows such results from a signal with measurements of vibrations from the NASA Bearing Dataset (https://www.kaggle.com/datasets/vinayak123tyagi/bearing-dataset) (1st test recorded 2003.11.25.23.39.56). As can be seen from the results, there are peaks in the energy of the wavelet representations near the temporal moments where there are also major peaks in signal.
Figure 11 shows corresponding results for an auditory percussion sound of a cymbal (from https://pixabay.com/sound-effects/search/cymbal/ item “Cymbal Crash”). As we can see, the time-causal wavelet transform largely responds over three ranges of temporal scales, with different durations of the responses over these ranges of temporal scales. In this way, the proposed time-causal wavelet transform leads to an intuitive separation of the temporal variations in the signal over different temporal scales.
@@ 8 Summary and discussion
Computationally, the proposed bandpass wavelet representations specifically allow from much more straightforward reconstruction of the original signal from the wavelet transforms. Whereas a standard (usually non-causal) wavelet representation requires reconstruction from the general inversion formula reproduced in Equation (7), when using the bandpass wavelet representation according to Equation (97), the original signal can instead be recovered from a mere summation of the bandpass wavelet channels according to Equation (101). Notably, this could also open up to real-time manipulation of local contents in the signal, by restricting the reconstruction to subsets of temporal subbands, that is over subsets of the temporal scale channels.
As we have shown in Equation (112), there is a direct relationship between the underlying computationally equivalent difference-of-time-causal-limit-kernels (DoT) kernel and the first-order temporal derivative of the time-causal limit kernel. This relationship is specifically qualitatively different from the corresponding relationship for non-causal wavelets based on derivatives of the Gaussian kernel, where the corresponding bandpass kernel does according to Equation (110) instead correspond to the second-order derivative of the non-causal Gaussian kernel.
Then, we have in Section 4 turned to the problem of numerically approximating the time-causal limit kernel and its temporal derivatives for discrete signals. In this regard, first-order recursive filters constitute the natural analogue to the first-order integrators that equivalently describe the temporal smoothing with truncated exponential kernels, and which constitute the main computational primitives in the continuous theory. Thereby, we have describe a canonical way to compute discrete time-causal wavelet representations in Section 5, and also shown how this representation relates to bandpass representations, as can also be defined from continuous wavelets. Notably, the proposed implementation method is fully time-recursive, and does therefore lend itself directly to real-time processing on regular signal processing architectures. Specifically, as described in Algorithm 1, the temporal smoothing can be implemented in terms of a low number (4-8) first-order recursive filters coupled in cascade. The wavelet transforms are then obtained by applying compact temporal difference operators according to Equation (123) to the temporally smoothed data, complemented by complementary scale normalization. Thereby, real-time computations can be achieved on very low-capacity computational architectures.
@@
Based on these results, and the above mentioned principled theoretical foundations, we propose that this notion of time-causal wavelet analysis could be a valuable tool for signal processing tasks, where streams of signals are to be processed and analyzed in real time. Beyond real-time monitoring and time-series analysis and prediction, such applications may concern purposes of real-time feedback for e.g. control-loop systems, and as time-causal and scale-covariant temporal basis functions in learning-based approaches. Specifically, we argue that the proposed time-causal wavelet representations could be suitable for signals that may contain local variations over a rich span of temporal scales, and more generally for analysing physical or biophysical temporal phenomena, where a fully time-causal analysis is called for to be physically realistic.
@@
Concerning further applications of the theory, we have intentionally in this presentation focused solely on theoretical and generic properties of the proposed families of time-causal wavelet representations, and not here aimed at developing more application-oriented experiments beyond the proof-of-concept studies in Sections 7 and 6.2. The motivation for this is that we want the results to be as timeless as possible, and not depending on specific properties of the experimental data, which may be different for different application domains, and may furthermore warrant the inclusion of additional more specialized computational mechanisms as well as integration with other signal processing approaches to build full-fledge applications.
Let us, however, emphasize that temporal derivatives of the time-causal limit kernel, as used as the mother wavelets in the here proposed time-causal wavelet representations, have been used for handling the temporal domain in models for spatio-temporal receptive fields for purposes of video analysis in classical computer vision (Jansson and Lindeberg 2018, Lindeberg 2018a; 2018b). Those results clearly demonstrate that the numerical properties of the temporal derivatives of the time-causal limit kernel can be used for computing highly useful results on real-world temporal data.
Temporal derivatives of the time-causal limit kernel have also been used for modelling the temporal component in biological visual receptive fields (Lindeberg 2021), and leading to provable covariance properties under geometric image transformations (Lindeberg 2025a). A precursor to the time-causal limit kernel, in terms of finite sets of truncated exponential kernels coupled in cascade, with logarithmically distributed temporal scale levels, has furthermore been used for modelling the temporal component in auditory receptive fields (Lindeberg and Friberg 2015). In this work, we instead make use of the temporal derivatives of the time-casual limit kernel for developing a time-causal wavelet theory, and do also describe structural relations to non-causal wavelet representations based on derivatives of the Gaussian kernel.
Closely related work of using single truncated exponential kernels for modelling neural computations in biological perception. memory processes and temporal learning has in turn been performed by Jacques (2022), Holt et al. (2022), Howard et al. (2023), Howard (2024) and Howard et al. (2025), with close relations to earlier work on temporal scale-space representations by Lindeberg and Fagerström (1996). Compared to the zero-order truncated exponential functions used as temporal basis functions for learning by Howard et al. and Holt et al., the time-causal wavelets in this paper obey simplifying cascade properties over temporal scales and are also DC-balanced. For biological support of the logarithmic distribution of the temporal scale levels used both here, according to Equation (12), and in those works, see Cao et al. (2022).
@@
The raw zero-order (non-differentiated) time-causal limit kernel has additionally been successfully used as a replacement of the non-causal Gaussian window function in time-causal time-frequency analysis. For the corresponding notion of a time-causal and time-recursive analogue of Gabor filtering, corresponding to a time-causal analogue of the Morlet wavelet (Morlet et al. 1982a; 1982b), see Lindeberg (2025b). Compared to such time-causal time-frequency analysis, the time-causal wavelets developed in this paper require significantly less computations and lead to shorter temporal delays, with and with a better ability to handle rapid transients. For more general notions of a non-causal wavelets based on Hermite transforms, and which constitute extensions of the wavelets based on first- and second-order derivatives of the Gaussian kernel to higher orders of differentiation, see Markett (1993), Glaeske (2000) and Pandey and Phukan (2020).